what is the equation of a line that is perpendicular to the line y=-2/5x-5 and passes through the point 6,1
Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept

__________________________________________________________________________
equation of a line that is perpendicular to the line y=-2/5x-5 and passes through the point 6,1
New Line: slope = 5/2 (Negative reciprocal of -2/5)
y = (5/2)x + b
1 = (5/2)6 + b
-14 = b
y = (5/2)x -14